OpenStudy (anonymous):

Find the vertex of the quadratic equation: y = 2x2 + 8x − 4

jimthompson5910 (jim_thompson5910):

Do you know how to find the x coordinate of the vertex?

OpenStudy (anonymous):

i dont think

jimthompson5910 (jim_thompson5910):

If you have the general equation \[\Large ax^2+bx+c=0\] then the x coordinate of that vertex is \[\Large x = -\frac{b}{2a}\]

jimthompson5910 (jim_thompson5910):

In the case of \(\Large y = 2x^2+8x-4\), we can see that a = 2, b = 8 and c = -4

jimthompson5910 (jim_thompson5910):

Does that help?

OpenStudy (anonymous):

yeah kinda

jimthompson5910 (jim_thompson5910):

So what do you get for the x coordinate of the vertex?

OpenStudy (anonymous):

i got -2

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

Since x = -2 is the x coordinate of the vertex, you can find the y coordinate of the vertex by plugging in this x value. So... \(\Large y = 2x^2+8x-4\) \(\Large y = 2(-2)^2+8(-2)-4\) I'll let you finish

OpenStudy (anonymous):

okay i got (-2,-12)

jimthompson5910 (jim_thompson5910):

you nailed it

OpenStudy (anonymous):

thank you for your help

jimthompson5910 (jim_thompson5910):

you're welcome